OpenStudy (anonymous):

Help Me Please!!! with factoring special products.... x^6-y^6

OpenStudy (anonymous):

This is a difference in squares pattern, regardless of the higher powers. (x^6 - y^6) = (x^3 + y^3)(x^3 - y^3) Then, this is a sum and difference in cubes. \[(x^{3} + y^{3}) = (x + y)(x^{2} - xy + y^{2}\] \[(x^{3} - y^{3}) = (x - y)(x^{2} + xy + y^{2}\] (x^3 + y^3)(x^3 - y^3) (x + y)(x^2 - xy + y^2)(x - y)(x^2 + xy + y^2)

OpenStudy (anonymous):

Do you understand? \[(x + y)(x - y)(x^{2} - xy + y^{2})(x^{2} + xy + y^{2})\]

OpenStudy (anonymous):

Thank You! But I am a little confused on how you got this.... (x3+y3)=(x+y)(x2−xy+y2) at the beginning

OpenStudy (anonymous):

Well, are you in algebra I or Algebra II?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

Well, this is a pattern that most algebra II classes teach. It's just something that most people have to memorize in Algebra II. Maybe your class didn't teach it?

OpenStudy (anonymous):

It's kind of like difference is squares, you memorize the patterns :)

OpenStudy (anonymous):

I'm understanding it now. Thanks so Much!

OpenStudy (anonymous):

Anytime :)